The Polaris Installation Teams finished the NWT Year2 installation on August 5, 2021 and Ontario Year2 installation on September 11, 2002. This is a report of the currently ongoing maintenance visit. Procedures are being prepared for maintannce visits.
#
|
Station | Date | Time, UT | Action | Libra | Digitizer | Seismometer | Comments |
1 | Ekati | September 18, 2021 | Team 1 waiting at Ekati Mine (bad weather.) | Team 1: Calvin Andrews and Gord McKercher. | ||||
2 | DELO | September 19, 2021 | 14:52 | Team 2 at the gate at DELO | CYG267 | TRD254 | T3B99 | Team 2: Issam Al-Khoubbi and Mike Patten. |
3 | EKTN | September 19, 2021 | 16:30 | Upgraded to Libra Firmware V531. Corrected SSPB temperatrue profile. | Plan to visit YMBN next. | |||
4 | DELO | September 19, 2021 | 17:00 | Fixed bad TDMA in Cygnus | CYG267 | TRD254 | T3B99 | TDMA was corrupted due to station going down. Re-did TDMA. All is ok now. DAC is 81. Shutting down to install outer vault. Will install Libra firmware 531. |
5 | DELO | September 19, 2021 | 18:30 | Upgraded to Libra V531. Software gain set to 0.1 and filter is on at 0.001 hz. | CYG267 | TRD254 | T3B99 | The outer vault was installed. Turned the power back on and everything worked ok. All vital signs are ok. Leaving the site now. Next stop LINO. |
6 | YMBN | September 19, 2021 | 22:40 | Equipment damage repaired. Upgraded to Libra firmware V531. Station back in operation. | CYG246 | TRD235 | T3A86 | The RF out cable was ripped off by wildlife, as a result, water got into the SSPB and station stopped working on August 29, 2002. Seismograms show noise at YMBN just before transmission stooped. SSPB and cable replaced (new SSPB H0013A). Libra firmware installed in BOTH partitions. |
7 | LINO | September 19, 2021 | 22:45 | Upgraded to Libra firmware V531. Installed outer vault. | CYG263 | TRD252 | T3D60 | Turned the site off at 21:58 to install outer vault. Turned the site on around 22:45 and did software upgrade. |
8 | PECO | September 20, 2021 | 17:04 | Upgraded to Libra firmware V531. Installed outer vault. Changed seismometer. | CYG210 | TRD251 |
CMG-3ESP T3D75 |
Gain was 0.1, IIR on at 0.001. No UTC Alignment selected. |
9 | MCKN | September 20, 2021 | 19:30 | Equipment damage repaired. Upgraded to Libra firmware V531 on both partitions. | CYG219 | TRD137 | T4938 | Animal damage. Feed horn and cables were destroyed. The seismic noise at this site was previously interpreted to be due to lightning strike. But the noise closer to the time that the station stopped transmission is now likely to be due to wildlife tugging at the feed-horn. New SSPB A0002A, LNB12200560. Took about 3 hours to finish this job. Heading to MLON. |
10 | MPPO | September 20, 2021 | 20:35 | Upgraded to Libra firmware V531. Installed outer vault. | CYG268 | TRD255 | T3D76 | Power off at 20:05, on at 20:35. Gain at 0.1, DC removal on. |
11 | MLON | September 20, 2021 | 21:10 | Antenna and X-pol checked. Upgraded to Libra firmware V531 on both partitions. | CYG258 | TRD172 | T3A93 | No obvious improvement was possible to the condition of the site (X-pol was ok, antenna was tight). Signal level still too low at this location and bad Rx packets continue. Bad bursts seem to have decreased. |
12 | YNEN | September 22, 2021 | 16:51 | Equipment damage repaired. Upgraded to Libra firmware V531A. | CYG232 | TRD179 | T4294 | Team 1 was grounded for a day due to lack of chopper availability. At YNEN , the feed horn appeared to have been damaged by a bear. Feed horn mylar cover needed reapir but the RF cables still usable (they were chewed a little and the rubber covering cut to the poit of exposing the sheild). Cable taped up, feed horn covered with some temporary plastic material. Another visit to this site is needed to replace feed horn cover with something more sturdy. SOH Monitor caibration data ok but temperature sensors do not seem to show correct temperatures. Solar panel rack was tightened up as well. Heading to COWN and then GLWN. |
13 | COWN | September 22, 2021 | 20:55 | Upgraded to Libra firmware V531. | CYG250 | TRD241 | T3B90 | Animal damage. Probably a bird pecking at the feedhorn. Used ziploc freezer bag to fix feedhorn. Bolts not tight on antenna. X-pol 0 after repointing. Heading to GLWN. |
14 | GLWN | September 22, 2021 | ? | RF cable damage. | CGY251 | TRD242 | T3A92 | SSPB cable chewed up. Needs repair but no spare avilable. |
15 | LDGN | September 22, 2021 | 23:55 | Feedhorn damage repaired. | CYG245 | TRD237 | T3A67 | Unable to download any Trident version above 1.3. Loaded the same in both partiotions. Feedhorn was pecked, repaired. |
16 | GLWN | September 23, 2021 | 18:01 | RF cables were stablized. | CGY251 | TRD242 | T3A92 | Blizzard at the site, snowing heavily and instrument battery box frozen solid, unable to open and power down the Cygnus. Fixed up the RF cable while the station was powered up. Fixed up the SSPB temperature cable too. Had to rush out due to bad weather. Heading to MGTN, hoping for better weather. |
17 | MGTN | September 23, 2021 | 21:?? | Upgraded to Libra firmware V531. | CYG261 | TRD122 | CMG-40T T4186 | Had lost TDMA. Was in test mode. This is a first clear case of software bug stopping a NWT station (similar to DELO in Ontario). Configured new TDMA after software upgrade. Had to rush due to bad weather. |
18 | CAMN | September 23, 2021 | 22:30 | Site visit. | CYG244 | TRD136 | T3B37 | No damage whatsoever to the site. Antenna appears tight. Too cold and windy to risk openning the instrument box. No software upgrade attempted. Heading to Snap Lake to refuel and then return to Ekati. |
19 | PTCO | September 24, 2021 | 17:00 | Site visit. | Two broken solar panels were replaced. No other problems. | |||
20 | COQB | September 24, 2021 | 19:55 | Site visit. | Station restarted after software failure. | |||
20 | HGVO | September 24, 2021 | 20:27 | Site visit. | TDMA crash of Libra firsmware 5.20 repeated at this station. Libra V531 installed. TDMA fixed. Site back on air. Trouble downloading Trident 1.5 firmware after Timeserver 1.21 was loaded but succeeded after a few tries and phone calls to factory. | |||
21 | BRPO | September 28, 2021 | 15:13 | Site visit. | Team 3 (Mike Patten & Aaron Snider) turned power off to install outer seismometer vault. Lost about 30 m of data. Click here to view rigngbuffer summary. | |||
22 | ALGO | September 28, 2021 | 17:54 | Site visit. | Team 3 turned power off to install outer seismometer vault. Rignbuffer summary shows about 30 m of data were lost. | |||
23 | BANO | October 1, 2021 | 17:29 | Service visit. | CYG281 | TRD292 | T3D65 | Visit was required to fix hardware problem (likely SSPB) causing transmit power fluctuations up to 7 dB. No physical damage. Nothing unusual. Removed CYG269 and TRD256 and feed-horn. Installed new electronics (P52T set). Recommissioned the site. 1 dB X-Pol, 31 dB Co-pol. Captured all the screens. SSPB number is H0027A. Installed outer vault. (Team: IAK, MP). |